OpenStudy (anonymous):

Convert the following into slope intercept form please 5. 3x + 5y = 8 6. -4x – 6y = 12 7. ½ x + y = 4 8.y – 2 = 3(x +1) 9. y – 5 = ½ (x – 4) 10. 6y – 3x = 2

OpenStudy (anonymous):

Remember that slope intercept form is y = mx + b So for problems 5, 6, 7, and 10, you need to get y by itself on the left and divide by whatever number is in front of it. ie) 5) 3x + 5y = 8 5y = 8 - 3x 5y = -3x + 8 y = -(3/5)x + 8/5 The slope is m = -3/5 and the y intercept is b =8/5 For problems 8 and 9, just move the extra numbers to the right!

OpenStudy (anonymous):

THANK YOU SO MUCH OH MY GOSH

OpenStudy (anonymous):

Welcome :)